# MEGA-Bench[Uncaptioned image]: Scaling Multimodal Evaluation to over 500 Real-World Tasks

### v1

@@ 4.4 Error Analysis
To understand the limitations of state-of-the-art VLMs, we analyze the GPT-4o (0513) results by manually identifying the error types over a subset of 255 tasks from the Core set. We use the CoT setting since the reasoning process helps determine the error type. Figure 7 (right) presents the error distribution. For GPT-4o, the lack of various reasoning capabilities (e.g., symbolic reasoning for planning/coding tasks, spatial or temporal reasoning for complex perception tasks, etc.) is the dominating failure mode on MEGA-Bench. Please refer to §E for the full definition of error types and detailed example-wise inspection results with different models.
@@ 5 Conclusion
Appendix A Details of Annotation Protocols
@@
A.1 The unified annotation format
@@
A.2 General task collection and creation guidelines
@@
In our annotation process, the first two task types are encouraged. The task reviewers strictly control the number of the third type and reject the task if an annotator submits many tasks of the third type. Appendix F shows the detailed data source of all tasks in MEGA-Bench.
@@
Documentation. When submitting the pull request, the annotator must write README documentation for each task. If the desired metric has not been implemented, the documentation should contain the specification described in the last point. Furthermore, the doc should record the data source (e.g., the Web, an existing dataset, etc.) and brief descriptions of the task. These descriptions are instrumental in helping the core contributors assign various keywords to the task and creating Appendix F to show the details of all tasks.
A.3 Tools for coordinating annotation and quality control
@@
The GitHub repository for task organization. We created a private GitHub repository for constructing MEGA-Bench. The repository’s main branch is protected, and all task submissions must go through pull requests (PRs). The core contributors serve as the task reviewers and discuss with task annotators in the pull request forum to ensure the task conforms to our data collection guidelines (§A.2). The code of our evaluation pipeline, including the model query and score computation, is maintained in the same repository. The core contributors submit pull requests to support different VLMs and add new evaluation metrics, and these PRs are cross-reviewed by other core contributors.
@@
Appendix B Taxonomy Tree and Multi-dimensional Keywords
@@
B.1 Details of the Taxonomy Structure
subsection B.1 shows the detailed structure of our application-driven task taxonomy. The first level defines the broad scope of use cases. At the second level, tasks are categorized into more specific domains. These first two levels guide the annotation process of our benchmark and are gradually updated/refined in the annotation process. The third level lists the concrete names of tasks or task groups. If the third-level node is a task group, the number of concrete tasks under this group is shown in the parenthesis.
@@ B.1 Details of the Taxonomy Structure
B.2 Statistics of each keyword dimension
@@
Appendix C Evaluation Details
@@ C.1 Prompt template
C.2 Model query details
@@
Claude-3.5-Sonnet (Anthropic, 2024) 64 8
@@
Pixtral 12B (Mistral, 2024) 48 6
@@ C.3 LLM-assisted metrics
C.4 Rule-based metrics
@@
C.5 Answer extraction from model response
@@
Appendix D Complete Multi-dimensional Breakdown Results
@@
D.1 Breakdown results on the skill dimension
@@
GPT-4o (0513) (OpenAI, 2024a) 63.4 55.1 68.0 61.2 44.2 56.3 22.9 58.2 39.4 61.7
Claude-3.5-Sonnet (Anthropic, 2024) 57.6 52.8 69.7 57.0 47.7 54.1 23.8 54.5 40.8 60.3
Gemini-1.5-Pro-002 (Google, 2024b) 57.5 51.4 69.8 55.1 42.6 52.0 23.9 54.7 38.5 49.9
Gemini-1.5-Flash-002 (Google, 2024b) 55.9 44.8 63.8 49.8 34.4 46.3 19.0 51.0 34.5 43.2
GPT-4O Mini (OpenAI, 2024b) 55.7 41.9 69.0 51.5 34.1 44.9 19.4 46.7 29.4 48.8
Qwen2-VL-72B (Alibaba, 2024) 56.8 46.3 60.5 53.4 37.8 49.8 22.0 50.9 35.1 53.9
InternVL2-Llama3-76B (Chen et al., 2024d) 52.6 33.3 57.8 43.4 29.8 38.2 17.0 42.7 29.5 41.0
LLaVA-OneVision-72B (Li et al., 2024a) 47.8 31.7 60.1 36.7 29.5 36.2 13.9 42.1 29.6 28.2
Qwen2-VL-7B (Alibaba, 2024) 49.4 33.3 52.2 40.1 28.2 37.1 14.7 41.1 27.6 40.0
Pixtral 12B (Mistral, 2024) 41.9 32.8 56.9 38.2 28.3 34.6 10.6 37.8 26.8 37.6
InternVL2-8B (Chen et al., 2024d) 39.7 27.1 47.0 31.9 24.1 28.2 8.3 32.6 23.2 27.9
Phi-3.5-Vision (Abdin et al., 2024) 36.8 24.1 46.7 28.6 21.7 25.5 8.9 30.5 21.5 24.7
MiniCPM-V2.6 (Yao et al., 2024) 40.7 23.7 48.8 29.9 18.3 26.0 8.7 31.8 19.7 24.8
LLaVA-OneVision-7B (Li et al., 2024a) 36.8 24.5 45.0 25.5 19.0 25.2 6.7 30.0 21.8 19.0
@@
D.2 Breakdown results on the input format dimension
@@
GPT-4o (0513) (OpenAI, 2024a) 47.8 56.4 50.0 55.9 49.1 60.5 53.2
Claude-3.5-Sonnet (Anthropic, 2024) 44.3 57.0 52.6 50.8 48.0 56.4 50.9
Gemini-1.5-Pro-002 (Google, 2024b) 42.9 55.8 48.7 55.0 42.9 45.9 50.3
Gemini-1.5-Flash-002 (Google, 2024b) 38.5 50.5 40.1 51.7 36.0 38.4 49.0
GPT-4O Mini (OpenAI, 2024b) 29.4 47.6 38.9 46.5 36.2 47.0 45.5
Qwen2-VL-72B (Alibaba, 2024) 36.2 50.8 42.1 49.6 42.9 53.6 49.9
InternVL2-Llama3-76B (Chen et al., 2024d) 28.7 45.0 34.7 42.7 31.4 36.2 39.6
LLaVA-OneVision-72B (Li et al., 2024a) 23.9 44.0 34.6 42.4 21.3 23.3 44.5
Qwen2-VL-7B (Alibaba, 2024) 26.2 34.8 32.2 40.5 29.0 38.1 41.1
Pixtral 12B (Mistral, 2024) 24.0 37.5 32.2 37.1 28.8 30.6 41.0
InternVL2-8B (Chen et al., 2024d) 10.9 29.4 28.0 33.8 20.1 22.8 34.8
Phi-3.5-Vision (Abdin et al., 2024) 15.4 27.9 26.1 34.1 17.5 18.6 24.7
MiniCPM-V2.6 (Yao et al., 2024) 7.6 31.0 21.6 31.6 18.6 21.2 35.3
@@
Llama-3.2-11B (Meta, 2024) 6.4 25.2 16.9 24.8 11.5 11.9 21.2
@@ V Videos
D.3 Breakdown results on the output format dimension
@@
GPT-4o (0513) (OpenAI, 2024a) 53.5 59.9 54.5 44.6 62.7 47.8
Claude-3.5-Sonnet (Anthropic, 2024) 50.4 52.8 54.6 44.9 58.4 49.3
Gemini-1.5-Pro-002 (Google, 2024b) 44.8 51.5 55.4 46.9 55.8 44.1
Gemini-1.5-Flash-002 (Google, 2024b) 38.7 44.8 47.8 37.0 54.5 39.6
GPT-4O Mini (OpenAI, 2024b) 41.2 44.2 39.9 36.3 57.1 38.9
Qwen2-VL-72B (Alibaba, 2024) 44.4 51.0 52.0 40.3 51.6 44.6
InternVL2-Llama3-76B (Chen et al., 2024d) 36.0 39.4 38.8 29.2 45.8 34.7
LLaVA-OneVision-72B (Li et al., 2024a) 28.6 37.1 39.9 30.7 42.9 25.9
Qwen2-VL-7B (Alibaba, 2024) 34.0 35.2 39.9 32.7 39.1 34.2
Pixtral 12B (Mistral, 2024) 30.7 36.4 30.1 32.1 41.7 31.8
InternVL2-8B (Chen et al., 2024d) 24.9 27.4 30.3 22.4 35.4 25.2
Phi-3.5-Vision (Abdin et al., 2024) 21.7 25.7 26.0 21.4 36.5 21.4
MiniCPM-V2.6 (Yao et al., 2024) 23.3 25.5 29.3 20.8 36.5 17.8
LLaVA-OneVision-7B (Li et al., 2024a) 20.2 25.4 28.0 22.0 31.3 18.3
Llama-3.2-11B (Meta, 2024) 12.3 15.8 19.3 15.0 30.0 16.4

### v2

@@ 4.4 Error Analysis
To understand the limitations of state-of-the-art VLMs, we analyze the GPT-4o (0513) results by manually identifying the error types over a subset of 255 tasks from the Core set. We use the CoT setting since the reasoning process helps determine the error type. Figure 7 (right) presents the error distribution. For GPT-4o, the lack of various reasoning capabilities (e.g., symbolic reasoning for planning/coding tasks, spatial or temporal reasoning for complex perception tasks, etc.) is the dominating failure mode on MEGA-Bench. Please refer to §F for the full definition of error types and detailed example-wise inspection results with different models.
@@ 5 Conclusion
Appendix A Single-image Setting: Results and Analyses
Table 2 in the main paper focuses on models with multi-image support. However, some open-source models are only trained with single images. To provide a feasible evaluation setting for these models, we create a single-image (SI) setting using the single-image tasks in MEGA-Bench, containing 273 and 42 tasks from the Core and Open-ended sets, respectively.
Evaluation setup. The Chain-of-Thought (CoT) prompting is used for Core SI tasks. To make the entire query contain only one image as needed by some single-image models, we drop the image input of the 1-shot demonstration example (“✗ demo im” column in the table). In this case, the 1-shot example only demonstrates the output format, which is necessary for inferring the correct answer. For those models already evaluated in Table 2, we calculate the task-level average scores on single-image tasks to obtain the “✓ demo im” results. Compared to Table 2, 3 single-image models are evaluated and added: Molmo-72B-0924 (Deitke et al., 2024), Molmo-7B-D-0924 (Deitke et al., 2024), and POINTS-Qwen2.5-7B (Liu et al., 2024b).
Evaluation results. Table 3 presents the evaluation results of the SI setting. The Core and Open-ended scores of the standard setting (with CoT prompting) are also in the table for reference. Some observations from the table are listed below:
$\bullet$ Single-image tasks are easier than multi-image tasks in general, and all models get higher scores in the SI setting than in the standard setting.
$\bullet$ GPT-4o has the best overall SI score, slightly higher than Claude 3.5 Sonnet (1022). Interestingly, GPT-4o mini overtakes Gemini-1.5-Flash-002 under the SI setting, suggesting that Gemini-1.5-Flash has pretty stable performance across different numbers of image inputs.
$\bullet$ NVLM-72B (Dai et al., 2024) has much better scores in the SI setting than in the standard setting, suggesting its training data might only contain single or a few images.
$\bullet$ Comparing the “✓ demo im” and “✗ demo im” results of open-source models, the image input in the 1-shot demonstration example is not well utilized by the models to better understand the task logic. Including the image input in the demonstration example makes the results much worse for models like Llama-3.2-11B.
More detailed breakdown results are available on our project page and the leaderboard (hosted with Hugging Face Spaces)
Table 3: The single-image (SI) setting results of MEGA-Bench. The Core set evaluation uses Chain-of-Thought (CoT) prompting. The “demo img” means the image input of the 1-shot demonstration example. The “✓ demo im” directly takes the single-image subset average from the full results in Table 2. The “✗ demo im” means the 1-shot demonstration example only demonstrates the output format, and the entire query has a single image. We report “✓ demo im” alone for the proprietary models because they have good multi-image support. For open-source models, we do additional evaluations with the “✗ demo im” setting and use it to compute the overall score.
Model Core
Core SI
Open
Open SI
Overall SI
✓ demo im ✗ demo im ✓ demo im ✗ demo im
GPT-4o (0513) (OpenAI, 2024a) 52.65 55.30 - 64.78 66.00 - 56.73
Claude-3.5-Sonnet (1022) (Anthropic, 2024b) 52.59 54.63 - 65.63 67.64 - 56.36
Claude-3.5-Sonnet (0620) (Anthropic, 2024a) 50.41 52.03 - 63.74 64.80 - 53.73
Gemini-1.5-Pro-002 (Google, 2024b) 48.22 49.14 - 58.58 58.15 - 50.34
GPT-4o mini (OpenAI, 2024b) 40.77 44.31 - 58.65 59.56 - 46.32
Gemini-1.5-Flash-002 (Google, 2024b) 41.89 43.48 - 56.91 57.87 - 45.40
Qwen2-VL-72B (Alibaba, 2024) 45.42 47.31 47.31 56.40 58.50 55.10 48.34
InternVL2-Llama3-76B (Chen et al., 2024d) 35.63 39.32 39.99 51.93 55.33 55.47 42.05
Molmo-72B-0924 (Deitke et al., 2024) - - 36.48 - - 44.66 37.58
NVLM-72B (Dai et al., 2024) 21.59 31.19 32.99 34.78 48.67 44.69 34.55
LLaVA-OneVision-72B (Li et al., 2024a) 29.74 31.77 31.26 45.99 46.12 44.26 32.99
Qwen2-VL-7B (Alibaba, 2024) 32.93 35.04 35.39 43.96 45.87 45.17 36.69
Pixtral-12B (Mistral, 2024) 31.36 34.87 34.37 45.66 44.03 44.17 35.68
Aria-MoE-25B (Li et al., 2024d) 28.90 31.67 31.79 51.03 50.92 51.37 34.40
InternVL2-8B (Chen et al., 2024d) 24.09 27.19 27.65 39.79 40.94 39.39 29.21
Phi-3.5-Vision-4B (Abdin et al., 2024) 23.00 25.72 25.61 39.48 44.61 42.72 27.89
POINTS-Qwen2.5-7B (Liu et al., 2024b) - - 25.51 - - 30.32 26.15
MiniCPM-V2.6-8B (Yao et al., 2024) 22.96 23.82 23.23 41.73 42.54 43.61 25.95
LLaVA-OneVision-7B (Li et al., 2024a) 21.36 22.70 23.68 33.98 36.44 38.71 25.69
Qwen2-VL-2B (Alibaba, 2024) 20.88 24.16 22.78 31.54 30.59 35.09 24.43
Molmo-7B-D (Deitke et al., 2024) - - 20.98 - - 35.70 22.95
Llama-3.2-11B (Meta, 2024) 16.00 17.34 20.79 31.73 34.29 38.61 23.17
Aquila-VL-2B-llava-qwen (Gu et al., 2024) 16.00 16.98 20.77 24.57 24.58 31.47 22.20
InternVL2-2B (Chen et al., 2024d) 13.14 13.83 12.07 23.86 24.28 28.52 14.26
Idefics3-8B-Llama3 (Laurençon et al., 2024) 8.96 9.13 8.94 32.11 33.25 32.31 12.06
Appendix B Details of Annotation Protocols
@@
B.1 The unified annotation format
@@
B.2 General task collection and creation guidelines
@@
In our annotation process, the first two task types are encouraged. The task reviewers strictly control the number of the third type and reject the task if an annotator submits many tasks of the third type. Appendix G shows the detailed data source of all tasks in MEGA-Bench.
@@
Documentation. When submitting the pull request, the annotator must write README documentation for each task. If the desired metric has not been implemented, the documentation should contain the specification described in the last point. Furthermore, the doc should record the data source (e.g., the Web, an existing dataset, etc.) and brief descriptions of the task. These descriptions are instrumental in helping the core contributors assign various keywords to the task and creating Appendix G to show the details of all tasks.
B.3 Tools for coordinating annotation and quality control
@@
The GitHub repository for task organization. We created a private GitHub repository for constructing MEGA-Bench. The repository’s main branch is protected, and all task submissions must go through pull requests (PRs). The core contributors serve as the task reviewers and discuss with task annotators in the pull request forum to ensure the task conforms to our data collection guidelines (§B.2). The code of our evaluation pipeline, including the model query and score computation, is maintained in the same repository. The core contributors submit pull requests to support different VLMs and add new evaluation metrics, and these PRs are cross-reviewed by other core contributors.
@@
Appendix C Taxonomy Tree and Multi-dimensional Keywords
@@
C.1 Details of the Taxonomy Structure
subsection C.1 shows the detailed structure of our application-driven task taxonomy. The first level defines the broad scope of use cases. At the second level, tasks are categorized into more specific domains. These first two levels guide the annotation process of our benchmark and are gradually updated/refined in the annotation process. The third level lists the concrete names of tasks or task groups. If the third-level node is a task group, the number of concrete tasks under this group is shown in the parenthesis.
@@ C.1 Details of the Taxonomy Structure
C.2 Statistics of each keyword dimension
@@
Appendix D Evaluation Details
@@ D.1 Prompt template
D.2 Model query details
@@
Claude-3.5-Sonnet (1022) (Anthropic, 2024a) 64 8
Claude-3.5-Sonnet (0620) (Anthropic, 2024a) 64 8
@@
Pixtral-12B (Mistral, 2024) 48 6
Aria-MoE-25B (Li et al., 2024d) 32 4
POINTS-Qwen2.5-7B (Liu et al., 2024b) 1 0
@@
Qwen2-VL-2B (Alibaba, 2024) 16 2
InternVL2-2B (Chen et al., 2024d) 18 2
Aquila-VL-2B-llava-qwen (Gu et al., 2024) 8 1
@@ D.3 LLM-assisted metrics
D.4 Rule-based metrics
@@
D.5 Answer extraction from model response
@@
Appendix E Complete Multi-dimensional Breakdown Results
@@
E.1 Breakdown results on the skill dimension
@@
Claude-3.5-Sonnet (1022) (Anthropic, 2024b) 59.1 54.9 65.7 60.8 48.9 56.9 29.1 55.1 43.2 62.2
GPT-4o (0513) (OpenAI, 2024a) 63.5 55.1 68.0 61.6 44.2 56.3 22.9 58.2 39.4 62.2
Claude-3.5-Sonnet (0620) (Anthropic, 2024a) 57.6 52.8 69.7 57.5 47.7 54.1 23.8 54.5 40.8 60.8
Gemini-1.5-Pro-002 (Google, 2024b) 57.5 51.4 69.8 55.3 42.6 52.0 23.9 54.7 38.5 50.2
Gemini-1.5-Flash-002 (Google, 2024b) 55.9 44.8 63.8 49.9 34.4 46.3 19.0 51.0 34.5 43.4
GPT-4o mini (OpenAI, 2024b) 55.7 41.9 69.0 51.7 34.1 44.9 19.4 46.7 29.4 49.0
Qwen2-VL-72B (Alibaba, 2024) 56.8 46.3 60.5 53.9 37.8 49.8 22.0 50.9 35.1 54.4
InternVL2-Llama3-76B (Chen et al., 2024d) 52.6 33.3 57.8 43.7 29.8 38.2 17.0 42.7 29.5 41.3
LLaVA-OneVision-72B (Li et al., 2024a) 47.8 31.7 60.1 36.7 29.5 36.2 13.9 42.1 29.6 28.3
NVLM-72B (Dai et al., 2024) 40.9 25.8 45.6 29.4 26.4 24.0 6.7 22.8 15.7 32.2
Qwen2-VL-7B (Alibaba, 2024) 49.4 33.3 52.2 40.3 28.2 37.1 14.7 41.1 27.6 40.2
Pixtral 12B (Mistral, 2024) 41.9 32.8 56.9 38.3 28.3 34.6 10.6 37.8 26.8 37.8
Aria-MoE-25B (Li et al., 2024d) 49.4 32.8 58.1 40.0 27.6 32.6 11.9 37.8 24.8 35.7
InternVL2-8B (Chen et al., 2024d) 39.7 27.1 47.0 32.0 24.1 28.2 8.3 32.6 23.2 28.1
Phi-3.5-Vision (Abdin et al., 2024) 36.8 24.1 46.7 28.7 21.7 25.5 8.9 30.5 21.5 24.8
MiniCPM-V2.6 (Yao et al., 2024) 40.7 23.7 48.8 30.0 18.3 26.0 8.7 31.8 19.7 25.0
LLaVA-OneVision-7B (Li et al., 2024a) 36.8 24.5 45.0 25.6 19.0 25.2 6.7 30.0 21.8 19.1
Qwen2-VL-2B (Alibaba, 2024) 31.3 20.8 41.4 25.7 17.6 22.2 6.2 26.5 17.3 23.7
@@
E.2 Breakdown results on the input format dimension
@@
Claude-3.5-Sonnet (1022) (Anthropic, 2024b) 44.2 55.6 55.6 54.3 48.9 60.5 49.5
GPT-4o (0513) (OpenAI, 2024a) 47.8 56.4 50.0 56.1 49.1 60.8 53.2
Claude-3.5-Sonnet (0620) (Anthropic, 2024a) 44.3 57.0 52.6 51.0 48.0 56.9 50.9
Gemini-1.5-Pro-002 (Google, 2024b) 42.9 55.8 48.7 55.0 42.9 46.3 50.3
Gemini-1.5-Flash-002 (Google, 2024b) 38.5 50.5 40.1 51.7 36.0 38.7 49.0
GPT-4o mini (OpenAI, 2024b) 29.4 47.6 38.9 46.5 36.2 47.2 45.5
Qwen2-VL-72B (Alibaba, 2024) 36.2 50.8 42.1 49.8 42.9 54.0 49.9
InternVL2-Llama3-76B (Chen et al., 2024d) 28.7 45.0 34.7 42.9 31.4 36.3 39.6
LLaVA-OneVision-72B (Li et al., 2024a) 23.9 44.0 34.6 42.5 21.3 23.4 44.5
NVLM-72B (Dai et al., 2024) 5.7 34.7 30.3 32.6 21.7 23.9 0.0
Qwen2-VL-7B (Alibaba, 2024) 26.2 34.8 32.2 40.7 29.0 38.2 41.1
Pixtral 12B (Mistral, 2024) 24.0 37.5 32.2 37.1 28.8 30.7 41.0
Aria-MoE-25B (Li et al., 2024d) 19.6 36.1 32.4 37.3 27.8 28.3 42.9
InternVL2-8B (Chen et al., 2024d) 10.9 29.4 28.0 33.9 20.1 22.8 34.8
Phi-3.5-Vision (Abdin et al., 2024) 15.4 27.9 26.1 34.1 17.5 18.7 24.7
MiniCPM-V2.6 (Yao et al., 2024) 7.6 31.0 21.6 31.8 18.6 21.2 35.3
@@
Qwen2-VL-2B (Alibaba, 2024) 13.4 24.9 19.6 28.8 16.3 19.1 25.2
Llama-3.2-11B (Meta, 2024) 6.4 25.2 16.9 24.9 11.5 11.9 21.2
Aquila-VL-2B-llava-qwen (Gu et al., 2024) 10.1 19.7 19.4 24.6 11.4 7.5 21.4
InternVL2-2B (Chen et al., 2024d) 11.9 14.9 16.3 20.1 10.5 5.7 19.0
@@ V Videos
E.3 Breakdown results on the output format dimension
@@
Claude-3.5-Sonnet (1022) (Anthropic, 2024b) 51.9 53.9 57.8 48.2 62.4 50.7
GPT-4o (0513) (OpenAI, 2024a) 53.9 59.9 54.5 44.6 62.7 48.0
Claude-3.5-Sonnet (0620) (Anthropic, 2024a) 50.7 52.8 54.6 44.9 58.4 49.7
Gemini-1.5-Pro-002 (Google, 2024b) 44.9 51.5 55.4 46.9 55.8 44.4
Gemini-1.5-Flash-002 (Google, 2024b) 38.7 44.8 47.8 37.0 54.5 39.9
GPT-4o mini (OpenAI, 2024b) 41.2 44.2 39.9 36.3 57.1 39.1
Qwen2-VL-72B (Alibaba, 2024) 44.7 51.0 52.0 40.3 51.6 45.0
InternVL2-Llama3-76B (Chen et al., 2024d) 36.3 39.4 38.8 29.2 45.8 34.8
LLaVA-OneVision-72B (Li et al., 2024a) 28.7 37.1 39.9 30.7 42.9 25.9
NVLM-72B (Dai et al., 2024) 22.9 27.9 18.5 23.3 32.2 27.9
Qwen2-VL-7B (Alibaba, 2024) 34.3 35.2 39.9 32.7 39.1 34.3
Pixtral 12B (Mistral, 2024) 30.8 36.4 30.1 32.1 41.7 31.9
Aria-MoE-25B (Li et al., 2024d) 30.9 29.3 32.8 30.9 45.2 30.4
InternVL2-8B (Chen et al., 2024d) 25.1 27.4 30.3 22.4 35.4 25.2
Phi-3.5-Vision (Abdin et al., 2024) 21.8 25.7 26.0 21.4 36.5 21.4
MiniCPM-V2.6 (Yao et al., 2024) 23.5 25.5 29.3 20.8 36.5 17.8
LLaVA-OneVision-7B (Li et al., 2024a) 20.3 25.4 28.0 22.0 31.3 18.3
Qwen2-VL-2B (Alibaba, 2024) 16.2 20.0 25.7 22.0 30.2 21.0
Llama-3.2-11B (Meta, 2024) 12.4 15.8 19.3 15.0 30.0 16.4
Aquila-VL-2B-llava-qwen (Gu et al., 2024) 11.9 18.5 22.1 19.9 23.3 12.3
InternVL2-2B (Chen et al., 2024d) 11.3 15.5 21.3 16.0 21.4 5.7